Assume 1,000 par and semiannual compounding

GHI Company bonds have a coupon rate of 3.63 percent, 4 years to maturity, and a current price of $1,086. What is the current yield on these bonds (in percent)? Answer to two decimals
Answer #1

Information provided:

Future value= $1,000

Time= 4 years*2= 8 semi-annual periods

Current price= present value= $1,086

Coupon rate= 3.63%/2= 1.8150% per semi-annual period

Coupon payment= 0.018159*1,000= $18.15

Current yield is calculated using the below formula:

Current Yield= Annual interest/Current price

                          = $18.15*2/ $1,086

                         = $36.30/ $1,086

                         = 0.0334*100 = 3.34%.

In case of any query, kindly comment on the solution.